OpenStudy (anonymous):

Can you guys help? I don't understand how you put it into a two-order system and how to solve it. A man invests $5,200, part at 4% and the balance at 3%. If his total income for the two investments is $194, how much money did he invest at each rate? $_______ at 4% and $________ 3%

OpenStudy (phi):

his total income means the sum of the interest from both investments do you know the formula to the amount of interest you get for P dollars ?

OpenStudy (anonymous):

No, that's literally all it gives me

OpenStudy (phi):

see http://www.wikihow.com/Calculate-Simple-Interest if they are asking this question, you must have seen this formula (in your notes?)

OpenStudy (phi):

I = P * r * t we have to assume t is 1 year ? r is the interest rate P is the amount invested

OpenStudy (phi):

we don't know how much was invested in each, but we know the dollars add up to 5200 call the amount invested at 3% x and the amount invested at 4% y x+y= 5200 the interest according to I= P * r * t (with t=1) is for x: x*0.03 for y: y * 0.04 the interest adds up to 194 0.03x + 0.04y = 194

OpenStudy (phi):

can you solve this system of equations? 0.03x + 0.04y = 194 x + y = 5200